Cheetahs running at top speed have been reported at an astounding 114km/h (about 71 mi/h) by observers driving alongside the animals. Imagine trying to measure a cheetah's speed by keeping your vehicle abreast of the animal while also glancing at your speedometer, which is registering 114km/h. You keep the vehicle a constant 8.0m from the cheetah, but the noise of the vehicle causes the cheetah to continously veer away from you along a circularpath of radius 92m. Thus, you travel along a circular path of radius 100m.

Cheetahs running at top speed have been reported a
Answer #1

V=114km/hr =31.66m/s

Time period of cheetah= 2*pi*r/v

=2*3.14*92/31.66 =18.24 sec

Centripetal accelaraton of cheetah = v^2/r =31.66^2/92 =10.89

Centripetal accelaraton of car = 31.66^2/100=10.02 m/s2